The perimeter of a rectangle is 52 inches. If the length of the rectangle is 2 inches more than twice its width, what are the demensions of the rectangle. Can you show me the operation to solve?
Answer by CharStar(110)

The formula for a rectangle is P=2(length+width)
W=x
L=2+w
52=2(w+w+2)
52=2w+2w+4 Subtract 4 from both sides and add the 2w + 2w
48=4W
12=W
Width = 12
Length = 14
Check you answer
p=2(12+14)
P=24+28
Perimeter is 52
I hope that helps you.